### v1

# From Structured Prompts to Open Narratives: Measuring Gender Bias in LLMs Through Open-Ended Storytelling

Abstract
Large Language Models (LLMs) have revolutionized natural language processing, yet concerns persist regarding their tendency to reflect or amplify social biases present in their training data. This study introduces a novel evaluation framework to uncover gender biases in LLMs, focusing on their occupational narratives. Unlike previous methods relying on structured scenarios or carefully crafted prompts, our approach leverages free-form storytelling to reveal biases embedded in the models. Systematic analyses show an overrepresentation of female characters across occupations in six widely used LLMs. Additionally, our findings reveal that LLM-generated occupational gender rankings align more closely with human stereotypes than actual labor statistics. These insights underscore the need for balanced mitigation strategies to ensure fairness while avoiding the reinforcement of new stereotypes.
1 Introduction
Large Language Models (LLMs) have rapidly become integral to natural language processing and decision support. Despite their capabilities, these models may inadvertently preserve or even amplify social biases rooted in training data. Accurately quantifying these biases is essential to ensure their fair and responsible deployment.
Previous studies have used various methods to assess LLMs’ gender biases in professional contexts. Some approaches involve structured decision-making tasks, such as choosing candidates from lists of names tied to different demographic groups (Nghiem et al., 2024; An et al., 2024). Other strategies test linguistic biases using prompts with ambiguous pronouns, e.g., “The doctor phoned the nurse because she was late” (Kotek et al., 2023). While effective, these methods are limited to predefined scenarios, risk oversimplifying biases, and may reveal the evaluation’s intent, allowing LLMs to adapt strategically.
To address these shortcomings, we propose a simple evaluation strategy that avoids predefined scenarios and conceals the intent of the test. Specifically, we request LLMs to generate a story opening and setting for a character in a specific occupation and analyze the gender cues in the output, such as names and pronouns (i.e., she or he). We execute the exact prompt multiple times for each occupation and evaluate the gender distribution across repeated outputs. This method eliminates the need for strictly designed scenarios or carefully tailored prompts, allowing biases to appear naturally within the context of free-form storytelling. It avoids explicitly directing the model’s attention toward gender-related considerations. Thus, it provides an authentic and comprehensive view of occupational gender stereotypes.
Based on this new evaluation strategy, we have the following observations. First, LLMs exhibit an apparent overrepresentation of female characters in their generated stories, with 58 of 100 occupations showing female dominance ($80\%$ or more characters are female) across the six LLMs tested. Interestingly, this female dominance is not limited to traditionally stereotypical female-associated occupations. A possible reason for this pattern is that LLM service providers may have intentionally adjusted the models to increase female representation, perhaps aiming to counteract historical gender biases in language data. This intentional weighting could lead to overrepresenting female characters in various occupations, not just those traditionally associated with women. Second, the rankings of occupations by gender ratios in LLM outputs align more closely with the perceived gender associations of occupations (as measured by human ratings) than the actual gender distributions in the labor force. This indicates that LLMs are more likely to mirror societal perceptions of gender roles than real-world demographic distributions, potentially amplifying biases rooted in societal stereotypes. Although this observation may seem contradictory to the first conclusion, it is not. The increased representation of females across all categories appears to lead many occupations to become female-dominated. However, when occupations are ranked by gender composition ratios, the resulting sequence still strongly aligns with human perceptions of gender associations.
The rest of the paper is organized as follows. In Section 2, we review previous work on quantifying the bias of LLMs. Section 3 introduces our methodology for evaluating LLM gender biases in occupations. Section 4 presents the discoveries based on our evaluation. Finally, we discuss and conclude the work in Section 5.
2 Related Work
Several studies have investigated the presence and impact of demographic biases in LLMs. These studies employ a range of methodologies and focus on different aspects of biases.
2.1 Evaluating Bias by LLMs’ Decisions on Strictly Designed Scenarios
Several studies have asked LLMs to make decisions or recommendations, allowing researchers to assess biases in these results (Nghiem et al., 2024; An et al., 2024; Salinas et al., 2023; Haim et al., 2024). For example, Nghiem et al. (2024) prompted LLMs to select the “best” candidates from lists of names associated with different racial and gender groups in 40 occupations. Their findings revealed a significant preference for white male names. An et al. (2024) similarly investigated the impact of applicant names on LLM-generated hiring decisions. They manipulated applicant names in the prompts to measure the influence of perceived ethnicity and gender on acceptance rates. Their results showed evidence of discrimination against Hispanic applicants, particularly males. Although these studies effectively demonstrate the presence of name-based biases, they are limited to strictly designed scenarios and may not be generalized to other occupational contexts or broader real-world applications.
2.2 Evaluating Bias by LLMs’ Word Choices on Carefully Designed Prompts
Another line of research has focused on analyzing the linguistic features of LLM-generated text to uncover biases (Wan et al., 2023; Kotek et al., 2023; Lee et al., 2024; Soundararajan and Delany, 2024; Kumar et al., 2024; Bas, 2024). For instance, Wan et al. (2023) analyzed gender biases in LLM-generated recommendation letters by examining both the language style and lexical content used to describe male and female candidates. They revealed significant gender bias in ChatGPT and Alpaca, with female candidates often described using more communal descriptions (e.g., kind, great to work with) while male candidates were described using more agentic descriptions (e.g., expert, a standout in the industry). Kotek et al. (2023) used ambiguous sentences with gendered occupations and pronouns to assess LLMs’ reliance on gender stereotypes and analyze their reasoning. They found that LLMs frequently rely on gender stereotypes to resolve pronouns, revealing an implicit bias in the models’ understanding of gender and occupations. And finding that LLMs often rationalize these biases with flawed logic. Lee et al. (2024) investigated the inherent biases and toxicity of LLMs in the Korean social context. They assign various personas to LLMs and evaluate the toxicity of the word choices. They found that specific persona-issue combinations could yield harmful content, indicating that LLMs can reflect and amplify social biases existing in particular cultural contexts.
Unlike previous approaches, we prompt LLMs to generate stories about a variety of professions and analyze the gender ratio of the characters in the stories. This method allows LLMs to describe open-ended stories rather than in restricted scenarios (e.g., Nghiem et al. (2024) and An et al. (2024)) and eliminates the need for meticulously designed prompt sentences (e.g., Kotek et al. (2023)). Thus, it has the advantage of capturing more natural and authentic biases embedded in the model’s narrative style, providing a broader and less constrained perspective on occupational gender stereotypes.
3 Evaluating Occupational Stereotypes
This section details four key stages of our methodology: (1) selecting occupations, (2) asking LLMs to generate story openings and settings, (3) classifying the gender of the main character in each story, and (4) analyzing the results for gender bias.
We selected Llama 3.2 3B (Meta AI, 2024), Gemma 2 (2B and 9B) (Google et al., 2024b), GPT-4o mini (OpenAI et al., 2024), Gemini 1.5 Flash and Gemini 1.5 Flash-8B (Google et al., 2024a) for evaluation because these models strike a balance between accessibility and representational capabilities, allowing us to conduct comprehensive evaluations without requiring complex infrastructure. Specifically, we build Llama 3.2 3B and Gemma 2 (2B and 9B) via Ollama, while the GPT-4o mini, Gemini 1.5 Flash, and Gemini 1.5 Flash-8B models were accessed through their official APIs, all models were evaluated using their default settings to reflect typical usage scenarios.
3.1 Occupation Selection
To create a list of professions for evaluation, we combined two sources of data: the U.S. Bureau of Labor Statistics (2023) (BLS), which reflects the current occupational landscape in the United States, and the gender stereotype rating study (GSR) (Kennison and Trofe, 2003), which represents human perceptions of gender stereotypes associated with occupations. The integration process involved matching professions across the two datasets. However, BLS data often grouped multiple related occupations into broader categories. For example, BLS defines “Writers and Authors” as a single category, while GSR assigns separate ratings for “Writer” and “Author.” To reconcile these differences, we map occupations from GSR to their closest corresponding categories in BLS, keeping the finer distinctions from GSR wherever possible.
The GSR dataset contains the participants’ ratings on perceived gender associations of various professions, measured on a scale of 1 to 7, where 1 represents the 0% male association and 7 represents 100% male association. We converted these ratings to a male association percentage:
Male Association $\text{Male Association Rate}=\frac{(\text{Rating}-1)}{6}\times 100\%$ (1)
3.2 Prompt Design and Story Generation
We ask LLMs to generate stories by the prompt:
Please provide a story opening and setting, including the full name of the main character, who is a [occupation].
For each profession, we initiated 75 independent conversation sessions with the LLMs, prompting them to generate one story per session. This approach ensures that each story is created in isolation, preventing any influence or carryover effects from other stories about the same profession.
3.3 Gender Classification of Story Characters
We extracted the character’s name from each story and utilized the nomquamgender package  (Van Buskirk et al., 2023) to classify the gender of the extracted names. We manually labeled the gender of the characters in 2000 stories to evaluate the accuracy of the package. We found that it achieved an accuracy of $98.35\%$, which we deemed sufficient for the following analyses.
3.4 Bias Analysis
For each occupation in a given LLM, we analyzed the gender distribution of characters in the generated stories, using it as an indicator of the model’s tendencies to associate genders with occupations.
To evaluate the gender stereotypes exhibited by each LLM, we compared the gender distribution of the story characters across occupations against two benchmarks: (1) the gender associations reported in  Kennison and Trofe (2003), which reflect human perceptions of gender stereotypes for professions; and (2) the actual gender distribution in the U.S. labor force reported by the U.S. Bureau of Labor Statistics (2023), representing real-world demographics. These comparisons help us assess whether the LLMs’ gender biases align with societal perceptions, real-world statistics, or diverge in ways that introduce new biases.
Figure 1: Occupation Gender Heatmap. The occupations are ranked from the most female-oriented(left) to the most male-oriented(right) according to the GSR.
Figure 2: The boxplot of the male ratio for all occupations generated by each LLM.
(a) LlaMA 3.2 3B
(b) Gemma 2 9B
(c) Gemma 2 2B
(d) GPT-4o mini
(e) Gemini 1.5 Flash
(f) Gemini Flash-8B
Figure 3: The male ratios of occupations in LLM-generated stories vs. real-world data from the U.S. BLS. The x-axis represents occupations, ordered from most female-oriented (left) to most male-oriented (right).
Table 1: Kendall’s Tau between LLMs and two benchmarks BLS and GSR. All results are significant as the $p$-values are all below 2.95e-14.
Model U.S. BLS GSR
Llama3.2 3B 0.6606 0.7041
Gemma2 9B 0.5929 0.6735
Gemma2 2B 0.6640 0.7347
Gpt-4o mini 0.5858 0.6765
Gemini 1.5 flash 0.5573 0.6252
Gemini 1.5 flash-8B 0.5471 0.6285
4 Discovery
Our experiments revealed notable findings on the behavior and biases of LLMs.
4.1 LLMs are Biased Toward Females in Occupational Narratives
Our first significant observation is the overrepresentation of female characters in a majority of occupations generated by LLMs. Figure 1 ranks occupations based on human ratings, from those perceived as the most female-oriented to those considered the most male-oriented. Darker blue indicates a higher proportion of female characters; darker red indicates a higher proportion of male characters. The heatmap shows that all tested LLMs demonstrate a significant tendency toward female characters. Figure 2 further illustrates the gender distribution across all occupations. The median gender ratios of real-world data (US BLS) and human evaluations are relatively balanced (with values of 0.473 and 0.565, respectively); human evaluations are skewed slightly toward males. In contrast, all LLMs exhibit a significant bias, with median male proportions falling below $20\%$, and in some cases, even below $10\%$. These findings highlight a significant disparity between societal realities and the outputs generated by these models, demonstrating both strong associations between specific occupations and genders and a marked bias in LLM-generated content.
Figure 3 compares the male ratios of occupations in LLM-generated stories with real-world data from the U.S. BLS. The results reveal that for a significant number of occupations, the male ratio in the generated stories falls below the real-world proportion. Among the 100 occupations analyzed, 58 featured female characters in $80\%$ or more of the generated stories, while only 5 occupations exhibited a similar dominance for male characters.
Interestingly, LLMs often depict neutral or slightly male-oriented jobs (e.g., lawyer, programmer) as female-dominated. This could be due to biases in the training data, such as too many examples of women in these roles, or because companies try too hard to avoid stereotypes and end up favoring female representation instead.
This phenomenon could have significant impacts. First, it might challenge stereotypes and promote broader gender representation in traditionally male-dominated fields. Second, it risks creating new stereotypes, oversimplifying social dynamics, or misrepresenting the composition of the workforce.
4.2 LLM Rankings Regarding Occupations’ Male Ratios Match Human Judgments
After generating occupational stories using the LLMs, we rank the professions by their male ratios in descending order. This ranking is compared to the corresponding rankings based on male-to-female ratios from the BLS data. Similarly, we compare the LLM-generated rankings to the rankings derived from human gender perception ratings. Table 1 presents the consistency of these rankings using Kendall’s Tau. The results indicate that the ranking consistency between human perceptions and LLM-generated rankings is higher than the consistency between BLS data and LLM-generated rankings. For instance, Llama 3.2 3B exhibited a Kendall’s Tau of $0.70$ ($p<0.001$) with human ratings, compared to $0.66$ ($p<0.001$) with labor statistics. Similarly, Gemma 2 2B demonstrated a Kendall’s Tau of $0.73$ ($p<0.001$) with human ratings, versus $0.66$ ($p<0.001$) with labor statistics.
We also find that smaller models, such as Llama 3.2 3B and Gemma 2 2B, align better with both human perceptions and real-world gender distributions compared to larger models like GPT-4o mini and Gemini 1.5 Flash. Thus, smaller models might retain a simpler relationship with their training data, leading to rankings that align more closely with external benchmarks. However, this trend is not universal across all cases (e.g., Gemma2 9B is better aligned than Gemini 1.5 flash-8B), indicating that the relationship between model size and bias may be complex and context-dependent. More research is needed to understand the underlying factors that contribute to this observation.
5 Discussion
Our study sheds light on gender biases in LLMs and their impact on real-world applications. This section explores key implications, possible causes, and fairness considerations in LLM deployment.
First, the overrepresentation of female characters across occupations in LLM-generated stories suggests that developers may attempt to correct historical gender imbalances in training data. Although such efforts are commendable and align with the goals of increasing inclusivity, they may unintentionally lead to overcorrection, which risks fostering new stereotypes, such as the notion that women dominate certain professions, regardless of real-world gender distributions. These unintended consequences highlight the need for strategies to achieve balance without distorting societal realities or reinforcing counterproductive biases.
Additionally, the disparity between human perception-based rankings and real-world labor statistics underscores the complex interplay between societal norms, training data, and model outputs. LLMs tend to mirror societal perceptions more closely than actual demographics, as evidenced by the alignment of model-generated rankings with human perception-based rankings. This raises critical questions about the role of training data and model fine-tuning in shaping the narratives LLMs produce and their potential to either challenge or reinforce existing stereotypes.
Our findings emphasize the ethical responsibilities of LLM developers to ensure fairness and inclusivity while minimizing unintended consequences. These responsibilities extend beyond training processes to include deployment strategies, user education, and ongoing monitoring of model outputs. For instance, users should be informed of the potential biases present in LLM-generated content and encouraged to critically evaluate outputs.
6 Limitation
There are some limitations to our work. First, it does not evaluate large-scale LLMs, limiting the generalizability of the findings. Second, real-world data from the US Bureau of Labor Statistics (BLS) focus only on the US, restricting cross-cultural applicability. Lastly, the analysis considers only binary gender categories (male and female), overlooking non-binary and gender-diverse identities.